# Leveraging (Biased) Information: Multi-armed Bandits with Offline Data

### v1

@@ C.3 Proof for instance dependent regret lower bound
Step (42) is by the definition of consistency. Step (43) is implied by our construction that in instance $P$, arm $a$ is sub-optimal with optiamlity gap $\Delta(a)\geq\epsilon\Delta(a)$, and in instance $Q$, arm $a$ is the unique optimal arm and other arms have optimality gap at least $\epsilon\Delta(a)$. Step (44) is by the Chain rule Theorem C.2, as well as the BH inequality. By our set up, we have
@@
C.4 Proof for instance independent regret lower bound
C.4.1 Proof for Theorem 4.10
@@ C.4 Proof for instance independent regret lower bound
$P^{\text{(on)}}_{a}=\begin{cases}{\cal N}(\Delta,1)&\quad\text{if }a=1\\ {\cal N}(0,1)&\quad\text{if }a\in{\cal A}\setminus\{1\}\end{cases}\text{, where }\Delta=\frac{1}{\sqrt{\tau_{*}}}.$
Consider the values $(\tilde{\tau},\tilde{n})$ defined as $\tilde{n}(a)=\mathbb{E}_{P,\pi}[N_{T}(a)]$ for each $a\in{\cal A}$, and $\tilde{\tau}=\min_{a\in{\cal A}}\{T_{\text{S}}(a)+\tilde{n}(a)\}$. Evidently, $(\tilde{\tau},\tilde{n})$ is feasible to (LP), and thus we have $\tilde{\tau}\leq\tau_{*}$, the optimum of (LP). In particular, there exists an arm $a\in{\cal A}$ such that $T_{\text{S}}(a)+\mathbb{E}_{P,\pi}[N_{T}(a)]\leq\tau_{*}$. To this end, let’s consider two situations:
@@
$Q^{\text{(on)}}_{a}=\begin{cases}{\cal N}(\Delta,1)&\quad\text{if }a=1\\ {\cal N}(2\Delta,1)&\quad\text{if }a=k\\ {\cal N}(0,1)&\quad\text{if }a\in{\cal A}\setminus\{1,k\}\end{cases}\text{, where }\Delta=\frac{1}{\sqrt{\tau_{*}}}.$
@@
$P^{\text{(off)}}_{a}=N(0,1)\text{ for all $a\in{\cal A}$, and }P^{\text{(on)}}_{a}=\begin{cases}{\cal N}(\Delta,1)&\quad\text{if }a=1,\\ {\cal N}(0,1)&\quad\text{if }a\in{\cal A}\setminus\{1\}.\end{cases}$
@@
$Q^{\text{(off)}}_{a}={\cal N}(0,1)\text{ for all $a\in{\cal A}$, and }Q^{\text{(on)}}_{a}=\begin{cases}{\cal N}(\Delta,1)&\quad\text{if }a=1,\\ {\cal N}(2\Delta,1)&\quad\text{if }a=a^{\prime},\\ {\cal N}(0,1)&\quad\text{if }a\in{\cal A}\setminus\{1,a^{\prime}\}.\end{cases}$
@@
Step (49) is again by Theorem C.1. Step (50) is by the Chain rule (Theorem C.2), and our observations on the constructed $P,Q$. Step (54) is by the choice of arm $a^{\prime}$ and the KL divergence between $P^{(\text{on})}_{a^{\prime}},Q^{(\text{on})}_{a^{\prime}}$. The strict inequality in (51) is by the case assumption that $2\sqrt{KT}>T\cdot(V_{\text{max}}+1/\tau_{*})>T\cdot V_{\text{max}}$, which implies $\frac{T}{K-1}\cdot\frac{V_{\text{max}}^{2}}{2}<\frac{K}{K-1}$. The larger than equal in (51) is by the assumption that $K\geq 2$. Altogether, the desire regret lower bound of
@@
largely by following Case 1b with a different $\Delta$, as well as the proof of Lemma 15.2 in (Lattimore & Szepesvári, 2020). Recall $\sqrt{KT}\geq T/\sqrt{\tau_{*}}$ since $\tau^{*}\geq T/K$, so the case condition implies $\sqrt{KT}\leq TV_{\text{max}}$, meaning $\sqrt{K/T}\leq V_{\text{max}}$. Now, set $\Delta=(1/2)\sqrt{(K-1)/T}$, and consider the Guassian instance with reward disribution $P$:
$P^{\text{(off)}}_{a}={\cal N}(0,1)\text{ for all $a\in{\cal A}$, and }P^{\text{(on)}}_{a}=\begin{cases}{\cal N}(\Delta,1)&\quad\text{if }a=1,\\ {\cal N}(0,1)&\quad\text{if }a\in{\cal A}\setminus\{1\}.\end{cases}$
@@
$Q^{\text{(off)}}_{a}={\cal N}(0,1)\text{ for all $a\in{\cal A}$, and }Q^{\text{(on)}}_{a}=\begin{cases}{\cal N}(\Delta,1)&\quad\text{if }a=1,\\ {\cal N}(2\Delta,1)&\quad\text{if }a=a^{\prime},\\ {\cal N}(0,1)&\quad\text{if }a\in{\cal A}\setminus\{1,a^{\prime}\}.\end{cases}$
@@
Step (52) is again by Theorem C.1. Step (53) is by the Chain rule (Theorem C.2), and our observations on the constructed $P,Q$. Step (54) is by the choice of arm $a^{\prime}$ and the KL divergence between $P^{(\text{on})}_{a^{\prime}},Q^{(\text{on})}_{a^{\prime}}$. Finally, putting in $\Delta=(1/2)\sqrt{(K-1)/T}$ gives us
@@
Appendix D More Details on Numerical Experiments
In all experiments, we set $\boldsymbol{\mu}^{\text{(on)}}=(0.8,0.5,0.5,0.5)$ (hence $a_{*}=1$). In Figure 1, we set $\boldsymbol{\mu}^{\text{(off)}}=(0.4,0.6,0.6,0.6)$, $V(1)=0.4$, and vary $V$ from 0.1 to 0.6. In Figure 2, we set $\boldsymbol{\mu}^{\text{(off)}}=(0.5,0.6,0.6,0.6)$, $V(1)=0.3$, and vary $T_{S}$ from 1000 to 3000, for $V=0.1,0.3$, respectively.

### v2

@@ D.2.1  Proof for Theorem 3
Step (52) is by the definition of consistency. Step (53) is implied by our construction that in instance $P$, arm $a$ is sub-optimal with optiamlity gap $\Delta(a)\geq\epsilon\Delta(a)$, and in instance $Q$, arm $a$ is the unique optimal arm and other arms have optimality gap at least $\epsilon\Delta(a)$. Step (54) is by the Chain rule Theorem 11, as well as the BH inequality. By our set up, we have
@@
D.3  Proof for instance independent regret lower bound for Multi-armed Bandits
D.3.1  Proof for Theorem 5
@@
$P^{\text{(on)}}_{a}=\begin{cases}{\cal N}(\Delta,1)&\qquad\text{if }a=1\\ {\cal N}(0,1)&\qquad\text{if }a\in{\cal A}\setminus\{1\}\end{cases}\text{, where }\Delta=\frac{1}{\sqrt{\tau_{*}}}.$
[…]
@@
$Q^{\text{(on)}}_{a}=\begin{cases}{\cal N}(\Delta,1)&\qquad\text{if }a=1\\ {\cal N}(2\Delta,1)&\qquad\text{if }a=k\\ {\cal N}(0,1)&\qquad\text{if }a\in{\cal A}\setminus\{1,k\}\end{cases}\text{, where }\Delta=\frac{1}{\sqrt{\tau_{*}}}.$
@@
$P^{\text{(off)}}_{a}=N(0,1)\text{ for all $a\in{\cal A}$, and }P^{\text{(on)}}_{a}=\begin{cases}{\cal N}(\Delta,1)&\qquad\text{if }a=1,\\ {\cal N}(0,1)&\qquad\text{if }a\in{\cal A}\setminus\{1\}.\end{cases}$
@@
[…]
@@
Step (59) is again by Theorem 10. Step (60) is by the Chain rule (Theorem 11), and our observations on the constructed $P,Q$. Step (64) is by the choice of arm $a^{\prime}$ and the KL divergence between $P^{(\text{on})}_{a^{\prime}},Q^{(\text{on})}_{a^{\prime}}$. The strict inequality in (61) is by the case assumption that $2\sqrt{KT}>T\cdot(V_{\text{max}}+1/\tau_{*})>T\cdot V_{\text{max}}$, which implies $\frac{T}{K-1}\cdot\frac{V_{\text{max}}^{2}}{2}<\frac{K}{K-1}$. The larger than equal in (61) is by the assumption that $K\geq 2$. Altogether, the desire regret lower bound of
@@
largely by following Case 1b with a different $\Delta$, as well as the proof of Lemma 15.2 in lattimore2020bandit. Recall $\sqrt{KT}\geq T/\sqrt{\tau_{*}}$ since $\tau^{*}\geq T/K$, so the case condition implies $\sqrt{KT}\leq TV_{\text{max}}$, meaning $\sqrt{K/T}\leq V_{\text{max}}$. Now, set $\Delta=(1/2)\sqrt{(K-1)/T}$, and consider the Guassian instance with reward disribution $P$:
$P^{\text{(off)}}_{a}={\cal N}(0,1)\text{ for all $a\in{\cal A}$, and }P^{\text{(on)}}_{a}=\begin{cases}{\cal N}(\Delta,1)&\qquad\text{if }a=1,\\ {\cal N}(0,1)&\qquad\text{if }a\in{\cal A}\setminus\{1\}.\end{cases}$
@@
[…]
@@
Step (62) is again by Theorem 10. Step (63) is by the Chain rule (Theorem 11), and our observations on the constructed $P,Q$. Step (64) is by the choice of arm $a^{\prime}$ and the KL divergence between $P^{(\text{on})}_{a^{\prime}},Q^{(\text{on})}_{a^{\prime}}$. Finally, putting in $\Delta=(1/2)\sqrt{(K-1)/T}$ gives us
@@
D.4  Proof for instance independent bound for Combinatorial Bandits with Linear Rewards
D.4.1  Proof for Theorem 8
The proof based on the special $m$-path bandit problem provided by kveton2015tight. Specifically, we define the $m$-path bandit problem as follows. The arm set is $\mathcal{A}=[K]$, where $K$ satisfies $K\geq 2m$ and $K/m$ is an integer. The feasible solution set $\mathcal{B}$ is defined as
$\mathcal{B}=\left\{\left\{\ell+(j-1)m:\ell=1,\cdots,m\right\}:j=1,\cdots\frac{K}{m}\right\}.$
In other words, the feasible solution set consists of $K/m$ paths, and each path contains $m$ arms. Both offline and online reward distributions satisfies the property that the distribution for the arms in the same path are identical. Specifically, $P^{\text{(off)}}_{\ell_{1}+(j-1)m}=P^{\text{(off)}}_{\ell_{2}+(j-1)m}$ and $P^{\text{(on)}}_{\ell_{1}+(j-1)m}=P^{\text{(on)}}_{\ell_{2}+(j-1)m}$ for any fixed $j$ and for any $\ell_{1},\ell_{2}\in[m]$. The offline data size $T_{\text{S}}(a)$ also satisfies that the size of dataset for the arms in the same path are the same, i.e. $T_{\text{S}}(\ell_{1}+(j-1)m)=T_{\text{S}}(\ell_{2}+(j-1)m)=T_{\text{S},j}$ for any fixed $j$ and for any $\ell_{1},\ell_{2}\in[m]$.
Notice that this problem is equivalent to a mult-armed bandit with offline data problem, with $K/m$ arms, each arm $j$ with $T_{\text{S},j}$ offline data, where $j\in[K/m]$, and the offline and online reward is scaled by $m$. Therefore, apply the result from Theorem 5, we obtain a lower bound
$\Omega\left(m\cdot\min\left\{\sqrt{\frac{K}{m}\cdot T},\left(\frac{1}{\sqrt{\tau_{*}^{\prime}}}+V_{\text{max}}\right)\cdot T\right\}\right)=\Omega\left(\min\left\{\sqrt{mKT},\left(\frac{1}{\sqrt{\tau_{*}^{\prime}}}+V_{\text{max}}\right)\cdot mT\right\}\right)$
where
$\tau_{*}^{\prime}=\max_{\tau,n}$ $\tau$
s.t. $\tau\leq T_{\text{S},j}+n_{j}\qquad\forall j\in[K/m],$
$\sum_{j\in[K/m]}n_{j}=T,$
$\tau\geq 0,n_{j}\geq 0\qquad\;\;\forall j\in[K/m].$
For each pair of feasible solution of the above optimization problem $(\tau,\{n_{j}\}_{j=1}^{K/m})$, the solution $(\tau,\{n(a)\}_{a\in\mathcal{A}})$, where
$n(\ell+(j-1)m)=n_{j},\ \forall\ell\in[m],\forall j.$
is a pair of feasible solution for the problem of (22) . Thus $\tau_{*}^{\prime}\leq\tau_{*}^{\text{C}}$, and the lower bound is proved.